OpenStudy (anonymous):

how to solve 4 2/5 tIMES 3.14 =??

OpenStudy (anonymous):

commone denominator. improper fraction of 4 with a denominator of 5 so 4 * (5/5) = 20/5 then 20/5 + 2/5 = 22/5 then 22/5 = 4.4 so 4.4 * 3.14 = 13.816

OpenStudy (anonymous):

ahh I see okay got it thank you... X]